OpenStudy (anonymous):

use the method of cylindrical shells to find the volume generated by rotating the region bounded by .... x=y^2 x=2 about y=-2

OpenStudy (dumbcow):

 the general shell method for rotating around horizontal axis is: \[V = 2 \pi \int\limits_{a}^{b}r(y) h(y) dy\] where r is radius and h is height from graph, r = 2+y h = 2-y^2 the bounds are where x=y^2 = 2 --> y = +-sqrt2 \[V = 2 \pi \int\limits_{-\sqrt{2}}^{\sqrt{2}}(2+y)(2-y^{2}) dy\]